# Game4Loc: A UAV Geo-Localization Benchmark from Game Data

### v1

@@ Introduction
Vision-based UAV geo-localization, as an independent onboard technology that can work independently of communication systems, enables UAVs to autonomously obtain GPS information even when GNSS communication fails. This UAV visual localization task could be refered as a special case of cross-view geo-localization (Deuser, Habel, and Oswald 2023; Zheng, Wei, and Yang 2020; Hu et al. 2018). Recent research formulates this as a cross-view image retrieval problem (Lin et al. 2022; Dai et al. 2023). Given a drone-view image, the goal is to retrieve a matching scene from a database of GPS-tagged satellite-view images to infer the current GPS information of the UAV. Compared to traditional hand-crafted feature extraction algorithms, deep learning based methods achieve higher accuracy and better generalization performance (Tian, Chen, and Shah 2017; Dusmanu et al. 2019). However, such superiority is built upon the training on a large amount of paired matching images from drone-view and satellite-view.
@@
In fact, some works already noticed the above problems, and attempted to address it from both task desgin and data construction perspectives. VIGOR (Zhu, Yang, and Chen 2021) introduces a beyond ont-to-one matching task for ground-satellite matching. DenseUAV (Dai et al. 2023) and UAV-VisLoc (Xu et al. 2024) are two continuous range real-world drone-satellite paired datasets. Both of them expand the retrieval task to localization; however, the former data construction method still does not align with practical scenarios, and the latter lacks a definition of data pair construction and task design. Additionally, these real-wolrd data are limited in terms of scenes, camera angles, and flight altitudes/attitudes, which restricts its generalization performance in diverse scenarios.
In light of the above problems, we propose aligning directly with practical tasks at the data construction level by expanding the original perfect matching to encompass partial matching as Fig. 1. Under our setting, the drone-satellite pairs are consturcted following the real-world scenarios, where drone-view images are retrieved from a gallery of satellite-view images containing partial matches. By consturcting such retrieval task, we can recreate the real-wolrd UAV visual geo-localization scenarios from the task design and evaluate the localization performance based on the retrieval results. Based on this, to replicate various drone flight conditions, we utilize commercial video games to simulate and collect a contiguous large-range of drone-satellite image pairs dataset GTA-UAV from multiple flight alittudes/attitudes, and various flight scenarios. In total, 33,763 drone-view images are collected from the entire game map, encompassing various scenes such as urban, mountain, desert, forest, field, and coast.
@@
  • We develop a weighted contrastive learning method weighted-InfoNCE to enable the model to learn this partial matching paradigm.
@@
Due to the comprehensive coverage of high-altitude reference data such as satellite and aerial imagery, most studies use GPS-tagged satellite imagery as the reference view for cross-view geolocalization. Among them, many datasets focus on the cross-view matching between ground-level and satellite-view (Lin, Belongie, and Hays 2013; Tian, Chen, and Shah 2017; Liu and Li 2019; Zhai et al. 2017; Zhu, Yang, and Chen 2021). Specifically, VIGOR (Zhu, Yang, and Chen 2021) doubts the perfect one-to-one matching data pairs and introduces the concept of beyond one-to-one retrieval in ground-satellite matching. University-1652 (Zheng, Wei, and Yang 2020) frist introduces the drone-view into the cross-view datasets, where each drone-satellite pair focuses on a target university building. Although the drone’s perspective can serve as a retrieval target, the task still not achieve geolocalization. In following works, DenseUAV (Dai et al. 2023) and SUES-200 (Zhu et al. 2023a) change discrete sampling into continuous sampling and consider different altitudes. Constrained by flight costs and the limitations of Google Earth simulation, the variety of shooting angles and altitudes reamins very limited. Most importantly, these datasets construction methods still adhere to the one-to-oen perfect matching paradigm and do not align with practical scenarios. UAV-VisLoc (Xu et al. 2024) is a recently released real high-altitude drone dataset where each drone-view image is geotagged, while no clear task desgin has been defined for this data yet.
@@
One of the first deep learning based geolocalization works by Workman et al. (Workman, Souvenir, and Jacobs 2015) demonstrates the superior accuracy and generalization of CNNs compared to traditional hand-crafted features. They simply utilize a L2 Loss to minimize the feature distance between cross-views and perform retrieval based on feature distances. Some works (Lin et al. 2015) adopt the idea of contrastive learning, reducing the distance between positive sample pairs. Vo et al. (Vo and Hays 2016) introduces a triplet loss, which brings positive samples closer to the anchor while pushing negative samples farther away. Further, Hu et al. includes a weight-shared NetVLAD-layer (Arandjelovic et al. 2016) to obtain better global descriptors. Yang et al. and Zhu et al. (Yang, Lu, and Zhu 2021; Zhu, Shah, and Chen 2022) explore the Transformer architecture in geolocalization to extract additional geometric properties. Specifically, Zhu et al.(Zhu et al. 2023b) proposes research on the unaligned case, i.e., the partial matching problem mentioned. However, their experiments are still conducted on aligned datasets. Sample4Geo (Deuser, Habel, and Oswald 2023) adopts the recent pre-training approach used in vision-language work CLIP (Radford et al. 2021), applying large batch size contrastive learning to cross-view data. They enhance the learning effect by constructing numerous hard negatives based on InfoNCE (van den Oord, Li, and Vinyals 2019).
@@
Given a filed of view (FOV) captured by the UAVs, our target is to construct a GPS-tagged reference satellite-view images set from a contiguous area and localize the drone by finding a matching field within it. Due to the varying flight altitudes and attitudes of UAVs, the FOV can cover multiple scales of the ground area. To accommodate the varying scales of drone-view, we divide the reference satellite-view of the entire coverage area into multiple hierachical tiles using a tiled map approach, where the ground resolution between different levels differing by a factor of two. Unlike the aligned one-to-one retrieval strong assumption of existing datasets in Tab. 1, we do not center-align the drone-satellite pairs. Instead, we use a collect-then-match approach, pairing them by calculating the overlapping of the ground area covered by the two views. In such arbitrarily sampling way, the relationship between pairs changes from perfectly aligned matching to partial matching. Refer to the definition of positive samples in VIGOR (Zhu, Yang, and Chen 2021), we attribute samples with a ground area intersection over union (IOU) greater than 0.39 as a positive pair, and IOU greater than 0.14 as a semi-positive pair. The positive pairs are considered as ground truth for retrieval for their highest match, while semi-positive pairs are complementary to the paritial matching learning. Such paritial matching, in contrast to the strong assumption of perfect matching, can be considered a more challenging retrieval task. On the basis of coarse retrieval, since each of our view data points is GPS-tagged, we can also evaluate the retrieval results at the distance level. This provides a foundation for fine localization in further research. Comparing to the existing datasets for UAV visual geo-localization as Tab. 1, our proposed GTA-UAV dataset offers higher flexibility and can cover a wider range of task scenrarios. We believe that our dataset complements existing UAV visual localization datasets and significantly bridge the gap between current research and practical applications.
@@
In light of the existing works (Richter et al. 2016; Ros et al. 2016; Kiefer, Ott, and Zell 2022) on synthetic data, we utilize Grand Thef Auto V (GTAV) as a simulation platform. We collect 33,763 drone-view images covering distinctive areas in the whole game map, including urban, mountain, desert, forest, field, and coast. To cover various flight altitudes and attitudes of UAVs, we simulate multiple flight heights ranging from $80m$ to $650m$, and multiple camera angle ranges for roll $\phi\in[-10^{\circ},10^{\circ}]$, pitch $\theta\in[-100^{\circ},-80^{\circ}]$ and yaw $\psi\in[-180^{\circ},180^{\circ}]$. The raw drone-view images are captured in $1920\times 1440$ with GPS tagged for meter-level evaluation. Based on the entire game map’s area of $81.3km^{2}$, we utilize a staellite map with a ground resolution of about $0.2m$ and divide it into a total of 8 hierarchical tiles. Each tile image has a pixel resolution of $256\times 256$, where the highest zoom level tiles having a ground resolutino of about $0.27m$. We collect totaling 14,640 tiles from zoom levels 4 to 7 as reference satellite-view set, to accommodate possible flight altitudes. For each drone-view image, we record the GPS information, flight altitude, flight attitude, and camera angle at the time of capture. By combining the FOV angle setting, we could approximate the ground area covered by the drone-view FOV. Then by enumerating the nearby satellite tiles from each level for each drone-veiw image, we set those with a ground coverage IOU greater than 0.39 as a positive drone-satellite pair, and the IOU between 0.14 and 0.39 as a semi-positive drone-satellite pair as shown in Fig. 2. The detailed construction process and dataset statistics are put in the supplementary.
@@
$\begin{split}\mathcal{L}_{\text{weighted-InfoNCE}}(F_{q},\alpha_{q},F_{R})=\\ &-\alpha_{q}\log\frac{\exp(F_{q}\cdot F_{r^{+}}/\tau)}{\sum_{i}^{R}\exp(F_{q}\cdot F_{r_{i}}/\tau)}\\ &-(1-\alpha_{q})\sum_{i}^{R}\log\frac{\exp(F_{q}\cdot F_{r_{i}^{+,-}}/\tau)}{\sum_{j}^{R}\exp(F_{q}\cdot F_{r_{j}}/\tau)}\\ &=\alpha_{q}\mathcal{L}_{\text{InfoNCE}}(F_{q},F_{R})+(1-\alpha_{q})\mathcal{L}_{\text{uniform-InfoNCE}}(F_{q},F_{R}),\end{split}$ (1)
@@
$\alpha_{q}=\sigma(k,\text{IOU}_{qr^{+}})=1-\frac{1}{1+\exp(-k\times\text{IOU}_{qr^{+}})},$ (2)
where $k$ is a hyper-parameter and higher value represents greater curvature change. When $k$ approaches infinity, the loss function degenerates into the standard InfoNCE. In one single batch with batch size $N$, there are $N$ positive/semi-positive paired samples with positive weights, and the reamining $N\times(N-1)$ combinations are regarded as negative samples. The loss function uses dot-production as the similarity measurement, where positive/semi-positive samples are pushed towards higher values and negative samples towards lower. Building on the original InfoNCE, we incorporate weights for positive/semi-positive sample pairs into the loss function, introducing a degree of flexibility. This allows the model to adapt the similarity loss based on the extend of partial matching.
@@
Triplet Loss ($\mathcal{L}_{\text{triplet}}$) 24.78% 46.99% 35.13% 58.79% 879.06m 46.55% 85.07% 62.95% 87.63% 252.88m
InfoNCE Loss ($\mathcal{L}_{\text{InfoNCE}}$) 35.83% 63.79% 48.08% 68.15% 576.41m 51.88% 89.75% 67.74% 88.85% 204.08m
InfoNCE Loss ($\mathcal{L}_{\text{InfoNCE}}$, w/. MES) 45.97% 71.43% 57.19% 71.48% 460.08m 65.89% 93.09% 77.84% 88.52% 193.19m
Ours ($\mathcal{L}_{\text{weighted-InfoNCE}}$, w/. MES) 55.91% 81.07% 66.56% 76.35% 342.05m 82.95% 97.86% 89.65% 90.48% 119.05m
@@
In our exeperiments the ViT-Base (Dosovitskiy et al. 2021) with patch-size $16\times 16$ and $64M$ parameters is adopted as the image encoding architecture. Both drone-view images and satellite-view images are resized to $384\times 384$ before feeding into the network. The hyper-parameter $k$ of weighted-InfoNCE is set to 5 as default, and the learnable temperature parameter $\tau$ is initialized to $1$. Following Sample4Geo (Deuser, Habel, and Oswald 2023), we employ Adam optimizer (Kingma and Ba 2017) with a initial learning rate of 0.0001 and a cosine learning rate scheduler to train each experiment for 10 epochs in batch size of 64. The flipping, rotation, and grid dropout are included as data augmentation for training. Both positive and semi-positive pairs are used for training by default if not specifically noted, and we conduct experiments on this in the subsequent subsections. The further details are put in the supplementary.
@@
For our GTA-UAV dataset, we compare the proposed method with previous SOTA training methods under both cross-area and same-area settings using positive $+$ semi-positive and positive-only as training data respectively. As results in Tab. 2, in the proposed paritial matching settings, our proposed weighted-InfoNCE achieves the best results across all metrics. Specifically, comparing to the previous SOTA method (Deuser, Habel, and Oswald 2023) using InfoNCE, our method improves the R@1 for 7.74%, and Dis@1 for 109.29m in the cross-area setting trained on positive $+$ semi-positive data. The results trained on positive $+$ semi-positive data have less retrieval accuracy comparing to the results only trained on positive data. This is because that the retrieval evaluation considers only the positive references as the correct result, which is precisely the training target of the positive data. However, for the localization task, the results trained on both positive and semi-positive data achieve better results in the SDM@3 and Dis@1 metrics. This is because the semi-positive data enable the model to learn a more comprehensive understanding of partial matching relationships. The further analysis of proposed weighted-InfoNCE are put in the supplementary.
@@
ConvNeXt-Base 80.39% 87.26% 89.13% 190.87m
Swinv2-B 78.27% 85.94% 88.70% 198.34m
ViT-Base/16 82.95% 89.65% 90.48% 119.05m
Table 6: Performance on GTA-UAV of different data scales.
[…]
Cross-Area
300 36.94% 61.12% 47.75% 61.64% 809.48m
3,000 48.43% 74.75% 59.98% 72.39% 482.25m
33,763 55.91% 81.07% 66.56% 76.35% 342.05m
79,852 56.01% 82.19% 66.93% 76.17% 343.10m
Same-Area
300 46.64% 73.13% 58.17% 71.68% 714.11m
3,000 65.56% 89.26% 75.95% 84.22% 368.58m
33,763 82.95% 97.86% 89.65% 90.48% 119.05m
79,852 84.45% 98.53% 90.71% 92.37% 76.81m
@@
In existing cross-view geo-localization (Deuser, Habel, and Oswald 2023; Hu et al. 2018; Toker et al. 2021; Zhu, Shah, and Chen 2022) research, CNNs and Transformers are widely explored for learning useful representations. Some studies make adaptive modifications to achieve better learning capabilities (Zhu, Shah, and Chen 2022; Hu et al. 2018; Zhu et al. 2023b). Unlike previous tasks, in the GTA-UAV cross-area task and its corresponding real-world scenarios, the generalization to unseen data in unkown scenes needs to be emphasized. Based on studies of model generalization (Hoyer, Dai, and Van Gool 2023; Ji et al. 2024) and SOTA geo-localization methods (Deuser, Habel, and Oswald 2023), we compare several standard architectures in Tab. 5. The results show that the ViT has the best performance under the same order of magnitude parameters.
Data Scale Evaluation
To further explore the data quality of the proposed GTA-UAV dataset, we validate the model performance under different training data scales. As shown in Tab. 6, under the same-area setting, the performance of models in retrieval and localization metrics imporve as the amount of data increasing. However, due to the limited game scenarios, the amount of effective data is also bounded. In the cross-area setting, which emphasizes generalization performance, the model performance stagnates despite the increase in data scale. Considering the effectiveness of the data, we select a quantity of 33,763 as the final dataset size.
@@
We evaluate different hyper-parameter value $k$ of proposed weighted InfoNCE in Tab. 7. Different values exhibit varying sensitivity to the positive weight, and all these results outperform when $k\to\infty$ (i.e., the standard InfoNCE).

### v2

@@ Introduction
Vision-based UAV geo-localization, as an independent onboard technology that can work independently of communication systems, enables UAVs to autonomously obtain GPS information even when GNSS communication fails. This UAV visual localization task could be refered as a special case of cross-view geo-localization (Deuser, Habel, and Oswald 2023; Zheng, Wei, and Yang 2020; Hu et al. 2018) and visual place recognition (Arandjelovic et al. 2016). Recent research formulates this as a cross-view image retrieval problem (Lin et al. 2022; Dai et al. 2023). Given a drone-view image, the goal is to retrieve a matching scene from a database of GPS-tagged satellite-view images to infer the current GPS information of the UAV. Compared to traditional hand-crafted feature extraction algorithms, deep learning based methods achieve higher accuracy and better generalization performance (Tian, Chen, and Shah 2017; Dusmanu et al. 2019). However, such superiority is built upon the training on a large amount of paired images from drone-view and satellite-view.
@@
In fact, some works already noticed the above problems, and attempted to address it from both task desgin and data construction perspectives. VIGOR (Zhu, Yang, and Chen 2021) introduces a beyond ont-to-one matching task for ground-satellite matching. DenseUAV (Dai et al. 2023) and UAV-VisLoc (Xu et al. 2024) are two continuous range real-world drone-satellite paired datasets. Both of them expand the retrieval task to localization; however, the former data construction method still does not align with practical scenarios, and the latter lacks a definition of data pair construction and task design. Additionally, these real-world data are limited in terms of scenes, camera angles, and flight altitudes/attitudes, which restricts its generalization performance in diverse scenarios.
In light of the above problems, we propose aligning directly with practical tasks at the data construction level by expanding the original perfect matching to encompass partial matching as Fig. 1. Under our setting, the drone-satellite pairs are consturcted following the real-world scenarios, where drone-view images are retrieved from a gallery of satellite-view images containing partial matches. By consturcting such retrieval task, we can recreate the real-world UAV visual geo-localization scenarios from the task design and evaluate the localization performance based on the retrieval results. Based on this, to replicate various drone flight conditions, we utilize commercial video games to simulate and collect a contiguous large-range of drone-satellite image pairs dataset GTA-UAV from multiple flight alittudes/attitudes, and various flight scenarios. In total, 33,763 drone-view images are collected from the entire game map, encompassing various scenes such as urban, mountain, desert, forest, field, and coast.
@@
  • We propose a weighted contrastive learning method weighted-InfoNCE to enable the model to learn this partial matching paradigm.
@@
Due to the comprehensive coverage of high-altitude reference data such as satellite and aerial imagery, most studies use GPS-tagged satellite imagery as the reference view for cross-view geolocalization. Among them, many datasets focus on the cross-view matching between ground-level and satellite-view (Lin, Belongie, and Hays 2013; Tian, Chen, and Shah 2017; Liu and Li 2019; Zhai et al. 2017; Zhu, Yang, and Chen 2021). Specifically, VIGOR (Zhu, Yang, and Chen 2021) doubts the perfect one-to-one matching data pairs and introduces the concept of beyond one-to-one retrieval in ground-satellite matching. University-1652 (Zheng, Wei, and Yang 2020) frist introduces the drone-view into the cross-view datasets, where each drone-satellite pair focuses on a target university building. Although the drone’s perspective can serve as a retrieval target, the task still not achieve geolocalization. In following works, DenseUAV (Dai et al. 2023) and SUES-200 (Zhu et al. 2023a) change discrete sampling into continuous sampling and consider different altitudes. Constrained by flight costs and the limitations of Google Earth simulation, the variety of shooting angles and altitudes reamins very limited. Most importantly, these datasets construction methods still adhere to the one-to-one perfect matching paradigm and do not align with practical scenarios. UAV-VisLoc (Xu et al. 2024) is a recently released real high-altitude drone dataset where each drone-view image is geotagged, while no clear task desgin has been defined for this data yet.
@@
One of the first deep learning based geolocalization works by Workman et al. (Workman, Souvenir, and Jacobs 2015) demonstrates the superior accuracy and generalization of CNNs compared to traditional hand-crafted features. They simply utilize a L2 Loss to minimize the feature distance between cross-views and perform retrieval based on feature distances. Some works (Lin et al. 2015; Vo and Hays 2016; Arandjelovic et al. 2016) adopt the idea of contrastive learning, reducing the distance between positive sample pairs. Yang et al. and Zhu et al. (Yang, Lu, and Zhu 2021; Zhu, Shah, and Chen 2022) explore the Transformer architecture in geolocalization to extract additional geometric properties. Specifically, Chen et al. (Chen et al. 2024) proposes research on the unaligned case, i.e., the non-centered or shifting targets. However, their experiments are still conducted on aligned datasets. Sample4Geo (Deuser, Habel, and Oswald 2023) adopts the recent pre-training approach used in vision-language work CLIP (Radford et al. 2021), applying large batch size contrastive learning to cross-view data. They enhance the learning effect by constructing numerous hard negatives based on InfoNCE (van den Oord, Li, and Vinyals 2019).
@@
Given a filed of view (FOV) captured by the UAVs, our target is to construct a GPS-tagged reference satellite-view images set from a contiguous area and localize the drone by finding a matching field within it. Due to the varying flight altitudes and attitudes of UAVs, the FOV can cover multiple scales of the ground area. To accommodate the varying scales of drone-view, we divide the reference satellite-view of the entire coverage area into multiple hierachical tiles, where the ground resolution between different levels differing by a factor of two. Unlike the aligned one-to-one retrieval strong assumption of existing datasets in Tab. 1, we do not center-align the drone-satellite pairs. Instead, we use a collect-then-match approach, pairing them by calculating the overlapping of the ground area covered by the two views. In such arbitrarily sampling way, the relationship between pairs changes from perfectly aligned matching to partial matching. Refer to the definition of positive samples in VIGOR (Zhu, Yang, and Chen 2021), we attribute samples with a ground area intersection over union (IOU) greater than 0.39 as a positive pair, and IOU greater than 0.14 as a semi-positive pair. The positive pairs are considered as ground truth for retrieval for their highest match, while semi-positive pairs are complementary to the paritial matching learning. Such paritial matching, in contrast to the strong assumption of perfect matching, can be considered a more challenging retrieval task. On the basis of coarse retrieval, since each of our view data points is GPS-tagged, we can also evaluate the retrieval results at the distance level. This provides a foundation for fine localization in further research. Comparing to the existing datasets for UAV visual geo-localization as Tab. 1, our proposed GTA-UAV dataset offers higher flexibility and can cover a wider range of task scenrarios. We believe that our dataset complements existing UAV visual localization datasets and significantly bridge the gap between current research and practical applications.
@@
In light of the existing works (Richter et al. 2016; Ros et al. 2016; Kiefer, Ott, and Zell 2022) on synthetic data, we utilize Grand Thef Auto V (GTAV) as a simulation platform. We collect 33,763 drone-view images covering distinctive areas in the whole game map, including urban, mountain, desert, forest, field, and coast. To cover various flight altitudes and attitudes of UAVs, we simulate multiple flight heights ranging from $80m$ to $650m$, and multiple camera angle ranges for roll $\phi\in[-10^{\circ},10^{\circ}]$, pitch $\theta\in[-100^{\circ},-80^{\circ}]$ and yaw $\psi\in[-180^{\circ},180^{\circ}]$. The raw drone-view images are captured in $1920\times 1440$ with GPS tagged for meter-level evaluation. Based on the entire game map’s area of $81.3km^{2}$, we utilize a staellite map with a ground resolution of about $0.2m$ and divide it into a total of 8 hierarchical tiles. Each tile image has a pixel resolution of $256\times 256$, where the highest zoom level tiles having a ground resolution of about $0.27m$. We collect totaling 14,640 tiles from zoom levels 4 to 7 as reference satellite-view set, to accommodate possible flight altitudes. For each drone-view image, we record the GPS information, flight altitude, flight attitude, and camera angle at the time of capture. By combining the FOV angle setting, we could approximate the ground area covered by the drone-view FOV. Then by enumerating the nearby satellite tiles from each level for each drone-veiw image, we set those with a ground coverage IOU greater than 0.39 as a positive drone-satellite pair, and the IOU between 0.14 and 0.39 as a semi-positive drone-satellite pair as shown in Fig. 2. The detailed construction process and dataset statistics are put in the supplementary.
@@
$\begin{split}\mathcal{L}_{\text{weighted-InfoNCE}}(F_{q},\alpha_{q},F_{R})=\\ &-\alpha_{q}\log\frac{\exp(F_{q}\cdot F_{r^{+}}/\tau)}{\sum_{i}^{R}\exp(F_{q}\cdot F_{r_{i}}/\tau)}\\ &-(1-\alpha_{q})\frac{1}{|R|}\sum_{i}^{R}\log\frac{\exp(F_{q}\cdot F_{r_{i}^{+,-}}/\tau)}{\sum_{j}^{R}\exp(F_{q}\cdot F_{r_{j}}/\tau)}\\ &=\alpha_{q}\mathcal{L}_{\text{InfoNCE}}(F_{q},F_{R})+(1-\alpha_{q})\mathcal{L}_{\text{uniform-InfoNCE}}(F_{q},F_{R}),\end{split}$ (1)
@@
$\alpha_{q}=\sigma(k,\text{IOU}_{qr^{+}})=\frac{1}{1+\exp(-k\times\text{IOU}_{qr^{+}})},$ (2)
where $k$ is a hyper-parameter and higher value represents greater curvature change. When $k$ approaches infinity, the loss function degenerates into the standard InfoNCE. In a single batch of size $N$, there are $N$ positive/semi-positive paired samples with corresponding positive weights from two views, and the loss function as Eq. 1 would be calculated twice symmetrically in two directions (drone to satellite, satellite to drone). The dot-production is utilized as the similarity measurement, where positive/semi-positive samples are pushed towards higher values. Building on the original InfoNCE, we incorporate weights for positive/semi-positive sample pairs into the loss function, introducing a degree of flexibility. This allows the model to adapt the similarity loss based on the extend of partial matching.
@@
Triplet Loss ($\mathcal{L}_{\text{triplet}}$) 24.78% 46.99% 35.13% 58.79% 879.06m 46.55% 85.07% 62.95% 83.63% 252.88m
InfoNCE Loss ($\mathcal{L}_{\text{InfoNCE}}$) 35.83% 63.79% 48.08% 68.15% 576.41m 52.67% 90.75% 67.74% 85.35% 204.08m
InfoNCE Loss ($\mathcal{L}_{\text{InfoNCE}}$, w/. MES) 45.97% 71.43% 57.19% 71.48% 460.08m 65.89% 93.09% 77.84% 86.52% 196.59m
Ours ($\mathcal{L}_{\text{weighted-InfoNCE}}$, w/. MES) 55.91% 81.07% 66.56% 76.35% 342.05m 84.95% 97.59% 90.15% 88.03% 149.07m
@@
In our exeperiments the ViT-Base (Dosovitskiy et al. 2021) with patch-size $16\times 16$ and $64M$ parameters is adopted as the image encoding architecture. Both drone-view images and satellite-view images are resized to $384\times 384$ before feeding into the network. The hyper-parameter $k$ of weighted-InfoNCE is set to 5 as default, and the learnable temperature parameter $\tau$ is initialized to $1$. Following Sample4Geo (Deuser, Habel, and Oswald 2023), we employ Adam optimizer (Kingma and Ba 2017) with a initial learning rate of 0.0001 and a cosine learning rate scheduler to train each experiment for 20 epochs in batch size of 64. The flipping, rotation, and grid dropout are included as data augmentation for training. Both positive and semi-positive pairs are used for training by default if not specifically noted, and we conduct experiments on this in the subsequent subsections. The further details are put in the supplementary.
@@
For our GTA-UAV dataset, we compare the proposed method with previous SOTA training methods under both cross-area and same-area settings using positive $+$ semi-positive and positive-only as training data respectively. As results in Tab. 2, in the proposed paritial matching settings, our proposed weighted-InfoNCE achieves the best results across all metrics. Specifically, comparing to the previous SOTA method (Deuser, Habel, and Oswald 2023) using InfoNCE, our method improves the R@1 for 20.08%, and Dis@1 for 234.36m in the cross-area setting trained on positive $+$ semi-positive data. The results trained on positive $+$ semi-positive data have less retrieval accuracy comparing to the results only trained on positive data. This is because that the retrieval evaluation considers only the positive references as the correct result, which is precisely the training target of the positive data. However, for the localization task, the results trained on both positive and semi-positive data achieve better results in the SDM@3 and Dis@1 metrics. This is because the semi-positive data enable the model to learn a more comprehensive understanding of partial matching relationships. The further analysis of proposed weighted-InfoNCE are put in the supplementary.
@@
ConvNeXt-Base 83.94% 89.54% 87.98% 160.49m
Swinv2-B 81.73% 88.32% 87.35% 196.06m
ViT-Base/16 84.95% 90.15% 88.03% 149.07m
@@
In existing cross-view geo-localization (Deuser, Habel, and Oswald 2023; Hu et al. 2018; Toker et al. 2021; Zhu, Shah, and Chen 2022) research, CNNs and Transformers are widely explored for learning useful representations. Some studies make adaptive modifications to achieve better learning capabilities (Zhu, Shah, and Chen 2022; Hu et al. 2018; Zhu et al. 2023b). Unlike previous tasks, in the GTA-UAV cross-area task and its corresponding real-world scenarios, the generalization to unseen data in unkown scenes needs to be emphasized. Based on studies of model generalization (Hoyer, Dai, and Van Gool 2023; Ji et al. 2024) and previous SOTA geo-localization methods (Deuser, Habel, and Oswald 2023; Zhu, Shah, and Chen 2022), we compare several standard architectures in Tab. 5. The results show that the ViT has the best performance under the same order of magnitude parameters. The practical commonly used architecture ResNet exhibits poor generalization ability, which may be attributed to its relatively weak representational and generalization capacities when dealing with significant variations in displacement, angles, and scenes. We also conduct experiments on the scale of model parameters in the supplementary.
@@
We evaluate different hyper-parameter value $k$ of proposed weighted InfoNCE in Tab. 6. There is a trade-off between treating partial matches as fully positive and maintaining flexibility (controlled by $k$), while all these results outperform when $k\to\infty$ (i.e., the standard InfoNCE). In addition, considering that the form of weighted-InfoNCE can be regarded as a weight-based label smoothing variant of InfoNCE, we also compare the results of InfoNCE with different fixed smooth value $\epsilon$ in the supplementary.